# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
___According to the Tenth Amendment, the Constitution lists all of the powers given to the United States government. Any power not listed in the Constitution belongs to either the American people or
Choices: [(a) the state governments, (b) the king or queen of Great Britain, (c) private businesses, (d) the Canadian government]
Answer index: 0
Subject: Social Studies
Skill: the-louisiana-purchase
@@
Description: In what year did the Louisiana Purchase take place?
Choices: [(a) 1792, (b) 1798, (c) 1803, (d) 1815]
Answer index: 2
Subject: Social Studies
Skill: the-lewis-and-clark-expedition
Description: Why did Thomas Jefferson and others want to find the Northwest Passage?
@@
Choices: [(a) It would make traveling less dangerous., (b) It would help lost explorers find their way home., (c) It would make trade with Asia easier.]
Answer index: 2
Subject: Social Studies
Skill: the-white-house
Description: Complete the sentence.
The White House is the home of ___.
@@
Choices: [(a) the president of the United States, (b) the governor of Maryland, (c) George Washington’s great-great-grandson, (d) the Speaker of the House of Representatives]
Answer index: 0
Subject: Social Studies
Skill: the-american-revolution-conclusion-and-aftermath
Description: Complete the sentence.
In 1779, ___ joined the Revolutionary War on the side of the United States.
Choices: [(a) Spain, (b) Canada, (c) Mexico, (d) Prussia]
Answer index: 0
Subject: Social Studies
Skill: the-american-revolution-the-boston-tea-party
Description: In 1770, colonists who wanted independence still had a hard time getting the colonies to work together. What made working together so difficult?
@@
Choices: [(a) The colonists were divided into two groups over the issue of slavery., (b) People living in different colonies often spoke different languages., (c) British soldiers would break up any large meetings of colonists., (d) The colonists were not well organized. They had no easy way to share information.]
Answer index: 3
Subject: Social Studies
Skill: the-constitution
Description: When was the Constitution written?
Choices: [(a) 1787, (b) 1867, (c) 1907, (d) 2007]
Answer index: 0
Subject: Social Studies
Skill: the-civil-war-war-tactics-and-the-home-front
Description: Complete the sentence.
@@
During the Civil War, the Union navy tried to close off Southern ports. The Union wanted to stop the Confederacy from trading with other countries. In other words, the Union wanted to create ___.
Choices: [(a) an ambush, (b) a blockade, (c) an armistice, (d) an assault]
Answer index: 1
Subject: Social Studies
Skill: the-great-depression-part-ii
Description: Where were the effects of the Great Depression felt?
@@
Choices: [(a) The Great Depression affected countries around the world., (b) The Great Depression only affected the United States., (c) The Great Depression’s impact was limited to North American countries., (d) The Great Depression was felt only in the Western Hemisphere.]
Answer index: 0
Subject: Social Studies
Skill: the-american-revolution-the-thirteen-colonies-under-british-rule
Description: How did Great Britain rule the Thirteen Colonies from so far away?
@@
Choices: [(a) Great Britain created a second national capital in the Thirteen Colonies., (b) Each colony elected leaders to visit the British government once a year., (c) Each colony had its own government, and those governments had to follow British law., (d) The king of Great Britain traveled to the Thirteen Colonies once a month.]
Answer index: 2
Subject: Social Studies
Skill: world-war-i-events-of-the-war
Description: Read the passage. Then answer the question.
The main type of fighting in World War I was trench warfare. Soldiers dug trenches, or deep holes in the ground, to protect themselves from enemy fire.
Soldiers stayed in the trenches for weeks at a time. The trenches were dirty and uncomfortable. Many soldiers got sick in the trenches.
Why were trenches used during World War I?
Choices: [(a) Trenches helped to end the war quickly., (b) Trenches protected soldiers from enemy weapons., (c) Trenches were clean places for soldiers to stay before returning to fight., (d) Trenches made "no man’s land" a safer place to be.]
Answer index: 1
Subject: Social Studies
Skill: antebellum-period-abolitionist-and-proslavery-perspectives
@@
Description: By 1860, there were almost four million enslaved African American people in the United States. What is an enslaved person?
Choices: [(a) a person who is drafted into the army, (b) a person who is owned by someone else, (c) a person who works for someone else for 4 to 7 years, (d) a person who is a criminal]
Answer index: 1
Subject: Social Studies
Skill: new-england-colonies-economy-and-conflict
Description: Slavery was not as widespread in New England as it was in the Southern Colonies. Why not?
@@
Choices: [(a) The colonists in New England had nothing valuable to trade for enslaved people., (b) The Puritans believed slavery was morally wrong. Most towns made slavery illegal., (c) New England did not have many large plantations, where enslaved labor was most valuable.]
Answer index: 2
Subject: Social Studies
Skill: jamestown-the-early-years
Description: Read the passage about the men who traveled to Jamestown. Then answer the question below.
Around 100 men signed up to travel to Jamestown. Here are some of the men who signed up:
However, around 50 of the men who signed up to go to weren’t workers at all! They were gentlemen. A gentleman is a man born into a noble family. Often, gentlemen are rich. They are rarely asked to do any work. Many gentlemen wanted to travel to Jamestown to find adventure.
Half of the first Jamestown settlers were gentlemen. Why might that have been a problem?
Choices: [(a) Gentlemen wanted to avoid adventures in the new colony., (b) Gentlemen did not have enough money to help start a new colony., (c) Gentlemen were not used to the hard work of starting a colony.]
Answer index: 2
Subject: Social Studies
Skill: middle-colonies-founding-and-government
Description: John Berkeley and Sir George Carteret wanted to make money by charging fees for land in New Jersey. But to do this, they needed to attract Europeans to settle the land.
Which of these rules would probably make Europeans more likely to move to New Jersey?
Choices: [(a) All settlers would have to work for seven years before they could get any land., (b) All settlers would have to pay taxes to the colonial government., (c) All settlers would be free to practice the religion of their choice.]
Answer index: 2
@@
Subject: Social Studies
Skill: antebellum-period-slavery-and-politics-part-ii
Description: While in prison in December 1859, John Brown handed a note to his guard. Use the text from that note to answer the question below.
I, John Brown, am now quite certain that the crimes of this guilty land will never be purged away; but with blood.
purged:removed
What did Brown believe would be needed to end slavery?
Choices: [(a) economic boycott, (b) petitions, (c) non-violent protest, (d) violence]
Answer index: 3
Subject: Social Studies
Skill: southern-colonies-economy-and-slavery
Description: Complete the sentence.
The largest farms in the Southern Colonies grew cash crops. Cash crops are plants meant ___.
Choices: [(a) to be sold, (b) to be given to the government, (c) to feed wildlife, (d) to be eaten by farmers]
Answer index: 0
Subject: Social Studies
Skill: presidential-elections
Description: Complete the sentence.
In an election, people choose their representatives and leaders by ___.
Choices: [(a) voting, (b) giving speeches, (c) drawing names randomly, (d) passing laws]
Answer index: 0
Subject: Social Studies
Skill: the-civil-war-the-end-of-the-war
Description: In 1864, the United States held a presidential election. Lincoln ran for re-election. A couple months before the election, Lincoln wrote the following note to his advisors. Use it to answer the question below.
@@
This morning, as for some days past, it seems exceedingly probable that this Administration will not be re-elected. Then it will be my duty to so cooperate with the President elect, as to save the Union between the election and the inauguration; as he will have secured his election on such ground that he can not possibly save it afterwards.
inauguration:when a president takes office
What did Abraham Lincoln think would happen if he lost the election?
Choices: [(a) The election would bring the country back together., (b) The new president would not be able to save the Union., (c) The new president would ask Lincoln to be vice president.]
Answer index: 1
Subject: Social Studies
Skill: mount-rushmore
@@
Description: When did workers start carving Mount Rushmore?
Choices: [(a) in 1863, (b) in 1888, (c) in 1901, (d) in 1927]
Answer index: 3
Subject: Social Studies
Skill: understand-overall-supply-and-demand
@@
Description: Eli Whitney invented the cotton gin in 1793. The cotton gin is a machine that makes it easier to harvest cotton quickly. What happened to the overall supply of cotton after 1793?
Choices: [(a) The supply went up., (b) The supply went down.]
Answer index: 0
Subject: Social Studies
Skill: costs-and-benefits
Description: Samuel is deciding whether to bake banana muffins or lemon muffins. He wants the muffins to be tasty. But he also wants to make a lot of muffins.
Suppose Samuel decides to bake lemon muffins. Which result would be a cost?
Choices: [(a) Samuel will get to eat more muffins. He can make more lemon muffins than banana muffins., (b) Samuel will give up the chance to eat banana muffins. He thinks banana muffins are tastier than lemon muffins.]
Answer index: 1
Subject: Social Studies
Skill: antebellum-period-economies-of-the-north-and-south
@@
Description: In the United States, most enslaved people worked on large farms that grew crops to sell. What were these farms called?
Choices: [(a) subsistence farms, (b) workshops, (c) factories, (d) plantations]
Answer index: 3
Subject: Social Studies
Skill: jamestown-growth-of-a-colony
@@
Description: In 1613, a Jamestown settler kidnapped Chief Powhatan’s favorite daughter. What was she called?
Choices: [(a) Pocahontas, (b) Massasoit, (c) Squanto, (d) Sacagawea]
Answer index: 0
Subject: Social Studies
Skill: world-war-ii-lead-up-to-war-in-europe-asia
Description: After the Treaty of Versailles was signed, a German newspaper stated the following. Read the passage. Then answer the question.
Today German honor is being carried to its grave. Do not forget it! The German people will, with unceasing labor, press forward to reconquer the place among the nations to which it is entitled. Then will come vengeance for the shame of 1919.
vengeance:payback
According to the passage, what does the author want Germany to do in the future?
Choices: [(a) The author wants Germany to join the Allied powers., (b) The author wants Germany to become a powerful nation again., (c) The author wants Germany to forget about World War I., (d) The author wants Germany to join the League of Nations.]
Answer index: 1
Subject: Social Studies
Skill: southern-colonies-founding-and-government
@@
Description: Many Virginians did not like the colonial government’s friendly policies toward Native Americans. In 1676, these Virginians violently rebelled against the government in Virginia. This event became known as Bacon’s Rebellion.
Which of the following is an example of a violent rebellion?
Choices: [(a) writing letters to the colonial governor of Virginia and the king, (b) voting in the next election for the Virginian colonial representatives, (c) marching on the colonial capital of Jamestown and burning it down]
Answer index: 2
Subject: Social Studies
Skill: middle-colonies-economy-and-society
Description: Merchants played a key role in trade between colonies and between the colonies and Europe.
Which of these things would a merchant be likely to do?
@@
Choices: [(a) use lumber from the Middle Colonies to build houses, (b) grow cash crops such as corn and wheat, (c) sell corn from the Middle Colonies to people in England, (d) make iron products such as plows and horseshoes]
Answer index: 2
Subject: Social Studies
Skill: world-war-ii-global-events
Description: Kathryn Mary Doody was a nurse at an army hospital in Hawaii when Pearl Harbor was bombed. Read Kathryn Mary Doody’s account of the events of December 7, 1941. Then answer the question.
@@
Pretty soon the night nurse came off duty, and she said to us, "Girls, you know what’s happening?" And we said, "No, what’s happening?" She said, "The island of Oahu has been attacked by the enemy, Japan." And I said, "Shut up." Because she was a person that always had the greatest stories to relate. And she said, "Don’t pay any attention to me. Go turn your radios on." So I went in the room and turned the radio on and that’s exactly what they were repeating, over and over again.. . .So I thought, "That means war. What’s war going to be like?"
According to the passage, what was Kathryn Mary Doody’s reaction when she first heard about the attack on Pearl Harbor?
Choices: [(a) She was scared and went to church to pray for peace., (b) She didn’t believe that it was true., (c) She told all of her friends that they should help the military., (d) She was angry at the United States government.]
Answer index: 1
Subject: Social Studies
Skill: early-mesopotamia
Description: Complete the text.
Mesopotamia is an area between the ___ Rivers. Because of the rich soil created by the two rivers, Mesopotamia is often called the Fertile Crescent.
Choices: [(a) Indus and Ganges, (b) Mississippi and Missouri, (c) Tigris and Euphrates, (d) Nile and Jordan]
Answer index: 2
@@
Subject: Social Studies
Skill: ancient-china
Description: For people in ancient Chinese communities, flooding around the Yellow and Yangtze Rivers was a common problem.
But floods had both positive and negative effects for people in ancient China. Which of the following was a positive effect of flooding?
@@
Choices: [(a) Floods killed many animals who lived near the river., (b) After a river flooded, fertile soil was left on the land., (c) Many crops were destroyed by the floods., (d) Floods destroyed homes located near the river.]
Answer index: 1
Subject: Social Studies
Skill: origins-of-islam
Description: Muslims believe there is only one god. They call this god Allah, which means "God" in Arabic. The following passage describes a Muslim belief about how Allah revealed, or communicated, messages to his followers. Read the passage. Then answer the question below.
We believe in Allah and [the message that] was revealed to us and what was revealed to Abraham . . . and the teachings which Allah gave to Moses and Jesus and to other prophets.
Al Imran 3:84 in Zafar Ishaq Ansari, translator, Towards Understanding the Qur’an, Volume 1. Copyright 1988 by The Islamic Foundation.
Based on this passage, which statement do most Muslims believe is true?
Choices: [(a) Allah is more powerful than the god Jews and Christians worship., (b) Allah is the same god that Jews and Christians worship., (c) Allah sent messages to Abraham, but not to other prophets.]
Answer index: 1
Subject: Social Studies
Skill: the-teachings-of-confucius
Description: Confucius’s teachings have been passed down in a text known as The Analects (AN-el-ekts). Read the passage about The Analects. Then answer the question below.
@@
The Analects is made up of discussions between Confucius and his students. When Confucius was alive, his students all remembered their own conversations with Confucius. When Confucius died, his students gathered together to write down their memories of Confucius’s teachings. The text they wrote down was called The Analects, which means "the selected speeches."
Complete the text.
collected and wrote down The Analects. The Analects is made up of ___.
Choices: [(a) Confucius . . . books Confucius wrote, (b) Confucius . . . conversations Confucius had with his students, (c) Confucius’s students . . . books Confucius wrote, (d) Confucius’s students . . . conversations Confucius had with his students]
Answer index: 3
Subject: Social Studies
Skill: ancient-egypt-the-old-kingdom
Description: Complete the text.
@@
Before Upper and Lower Egypt became one kingdom, they were divided into nomes, or ___. Each nome was controlled by a nomarch, or local king.
Choices: [(a) territories, (b) farms, (c) churches, (d) teams]
Answer index: 0
Subject: Social Studies
Skill: the-mauryan-empire
Description: One of Chandragupta’s most important military and political advisors was a man named Kautilya. Read the following story about Kautilya and Chandragupta. Then answer the question below.
Kautilya and Chandragupta gathered an army to overthrow the Nanda Empire. At first, their soldiers attacked the center of the Nanda Empire, but they were defeated by the powerful Nanda army.
According to tradition, Kautilya and Chandragupta took that mother’s advice about eating hot bread and used it to defeat the Nanda Empire.
How did Kautilya and Chandragupta use the mother’s advice to defeat the Nanda Empire?
@@
Choices: [(a) They convinced members of the Nanda army to change sides and help Chandragupta overthrow the empire., (b) They attacked the Nanda Empire from the edges, where it was weakest, rather than from the center where it was strong., (c) They attacked the center of the Nanda Empire until the army was too tired and weak to keep fighting.]
Answer index: 1
Subject: Social Studies
Skill: origins-of-christianity
Description: Complete the text.
Jesus of Nazareth is the most important figure in Christianity. Most Christians believe he ___.
Choices: [(a) is God, (b) built a temple in Jerusalem, (c) was a famous military leader, (d) was a Roman emperor]
Answer index: 0
Subject: Social Studies
Skill: ancient-egyptian-religion
Description: In ancient Egypt, powerful gods were an important part of religious beliefs. People who believe in many gods are called polytheists (pol-ee-THEE-ists).
Which statement would support the idea that ancient Egyptians were polytheists?
@@
Choices: [(a) Many ancient Egyptians worshipped the gods Osiris, Isis, and Horus., (b) Ancient Egyptian religious traditions lasted over 3,000 years., (c) Many different creation stories were popular in ancient Egypt., (d) Ancient Egypt was located in northeastern Africa, along the Nile River.]
Answer index: 0
Subject: Social Studies
Skill: the-gupta-empire
Description: The Gupta Empire was founded by a man who became known as Chandra Gupta (CHUN-druh GOOP-tuh). Read about Chandra Gupta. Then complete the sentence below.
Chandra Gupta was born to a ruling family, but they only controlled a small area and were not very powerful. That began to change around 320 CE when Chandra Gupta married Princess Kumaradevi.
Princess Kumaradevi came from a powerful family that controlled an area along the Ganges River. By marrying Kumaradevi, Chandra Gupta expanded the area he controlled, and he began forming an empire.
Complete the following sentence.
Chandra Gupta began the Gupta Empire after he ___.
@@
Choices: [(a) threatened to stop trading with regions that refused to let him rule, (b) inherited the large region his family had controlled for generations, (c) married a princess and took control of the area her family ruled, (d) won battles against many kingdoms along the Ganges River]
Answer index: 2
Subject: Social Studies
Skill: origins-of-judaism

### v2

@@ Appendix D Summary of Skills
Subject: Social Studies Skill: the-bill-of-rights Description: Complete the sentences. ___According to the Tenth Amendment, the Constitution lists all of the powers given to the United States government. Any power not listed in the Constitution belongs to either the American people or Choices: [(a) the state governments, (b) the king or queen of Great Britain, (c) private businesses, (d) the Canadian government] Answer index: 0
@@
Subject: Social Studies Skill: the-louisiana-purchase Description: In what year did the Louisiana Purchase take place? Choices: [(a) 1792, (b) 1798, (c) 1803, (d) 1815] Answer index: 2
@@
Subject: Social Studies Skill: the-lewis-and-clark-expedition Description: Why did Thomas Jefferson and others want to find the Northwest Passage? Choices: [(a) It would make traveling less dangerous., (b) It would help lost explorers find their way home., (c) It would make trade with Asia easier.] Answer index: 2
@@
Subject: Social Studies Skill: the-white-house Description: Complete the sentence. The White House is the home of ___. Choices: [(a) the president of the United States, (b) the governor of Maryland, (c) George Washington’s great-great-grandson, (d) the Speaker of the House of Representatives] Answer index: 0
Subject: Social Studies Skill: the-american-revolution-conclusion-and-aftermath Description: Complete the sentence. In 1779, ___ joined the Revolutionary War on the side of the United States. Choices: [(a) Spain, (b) Canada, (c) Mexico, (d) Prussia] Answer index: 0
@@
Subject: Social Studies Skill: the-american-revolution-the-boston-tea-party Description: In 1770, colonists who wanted independence still had a hard time getting the colonies to work together. What made working together so difficult? Choices: [(a) The colonists were divided into two groups over the issue of slavery., (b) People living in different colonies often spoke different languages., (c) British soldiers would break up any large meetings of colonists., (d) The colonists were not well organized. They had no easy way to share information.] Answer index: 3
Subject: Social Studies Skill: the-constitution Description: When was the Constitution written? Choices: [(a) 1787, (b) 1867, (c) 1907, (d) 2007] Answer index: 0
@@
Subject: Social Studies Skill: the-civil-war-war-tactics-and-the-home-front Description: Complete the sentence. During the Civil War, the Union navy tried to close off Southern ports. The Union wanted to stop the Confederacy from trading with other countries. In other words, the Union wanted to create ___. Choices: [(a) an ambush, (b) a blockade, (c) an armistice, (d) an assault] Answer index: 1
@@
Subject: Social Studies Skill: the-great-depression-part-ii Description: Where were the effects of the Great Depression felt? Choices: [(a) The Great Depression affected countries around the world., (b) The Great Depression only affected the United States., (c) The Great Depression’s impact was limited to North American countries., (d) The Great Depression was felt only in the Western Hemisphere.] Answer index: 0
@@
Subject: Social Studies Skill: the-american-revolution-the-thirteen-colonies-under-british-rule Description: How did Great Britain rule the Thirteen Colonies from so far away? Choices: [(a) Great Britain created a second national capital in the Thirteen Colonies., (b) Each colony elected leaders to visit the British government once a year., (c) Each colony had its own government, and those governments had to follow British law., (d) The king of Great Britain traveled to the Thirteen Colonies once a month.] Answer index: 2
Subject: Social Studies Skill: world-war-i-events-of-the-war Description: Read the passage. Then answer the question. The main type of fighting in World War I was trench warfare. Soldiers dug trenches, or deep holes in the ground, to protect themselves from enemy fire. Soldiers stayed in the trenches for weeks at a time. The trenches were dirty and uncomfortable. Many soldiers got sick in the trenches. Why were trenches used during World War I? Choices: [(a) Trenches helped to end the war quickly., (b) Trenches protected soldiers from enemy weapons., (c) Trenches were clean places for soldiers to stay before returning to fight., (d) Trenches made "no man’s land" a safer place to be.] Answer index: 1
@@
Subject: Social Studies Skill: antebellum-period-abolitionist-and-proslavery-perspectives Description: By 1860, there were almost four million enslaved African American people in the United States. What is an enslaved person? Choices: [(a) a person who is drafted into the army, (b) a person who is owned by someone else, (c) a person who works for someone else for 4 to 7 years, (d) a person who is a criminal] Answer index: 1
@@
Subject: Social Studies Skill: new-england-colonies-economy-and-conflict Description: Slavery was not as widespread in New England as it was in the Southern Colonies. Why not? Choices: [(a) The colonists in New England had nothing valuable to trade for enslaved people., (b) The Puritans believed slavery was morally wrong. Most towns made slavery illegal., (c) New England did not have many large plantations, where enslaved labor was most valuable.] Answer index: 2
Subject: Social Studies Skill: jamestown-the-early-years Description: Read the passage about the men who traveled to Jamestown. Then answer the question below. Around 100 men signed up to travel to Jamestown. Here are some of the men who signed up: However, around 50 of the men who signed up to go to weren’t workers at all! They were gentlemen. A gentleman is a man born into a noble family. Often, gentlemen are rich. They are rarely asked to do any work. Many gentlemen wanted to travel to Jamestown to find adventure. Half of the first Jamestown settlers were gentlemen. Why might that have been a problem? Choices: [(a) Gentlemen wanted to avoid adventures in the new colony., (b) Gentlemen did not have enough money to help start a new colony., (c) Gentlemen were not used to the hard work of starting a colony.] Answer index: 2
Subject: Social Studies Skill: middle-colonies-founding-and-government Description: John Berkeley and Sir George Carteret wanted to make money by charging fees for land in New Jersey. But to do this, they needed to attract Europeans to settle the land. Which of these rules would probably make Europeans more likely to move to New Jersey? Choices: [(a) All settlers would have to work for seven years before they could get any land., (b) All settlers would have to pay taxes to the colonial government., (c) All settlers would be free to practice the religion of their choice.] Answer index: 2
@@
Subject: Social Studies Skill: antebellum-period-slavery-and-politics-part-ii Description: While in prison in December 1859, John Brown handed a note to his guard. Use the text from that note to answer the question below. I, John Brown, am now quite certain that the crimes of this guilty land will never be purged away; but with blood. purged:removed What did Brown believe would be needed to end slavery? Choices: [(a) economic boycott, (b) petitions, (c) non-violent protest, (d) violence] Answer index: 3
Subject: Social Studies Skill: southern-colonies-economy-and-slavery Description: Complete the sentence. The largest farms in the Southern Colonies grew cash crops. Cash crops are plants meant ___. Choices: [(a) to be sold, (b) to be given to the government, (c) to feed wildlife, (d) to be eaten by farmers] Answer index: 0
Subject: Social Studies Skill: presidential-elections Description: Complete the sentence. In an election, people choose their representatives and leaders by ___. Choices: [(a) voting, (b) giving speeches, (c) drawing names randomly, (d) passing laws] Answer index: 0
@@
Subject: Social Studies Skill: the-civil-war-the-end-of-the-war Description: In 1864, the United States held a presidential election. Lincoln ran for re-election. A couple months before the election, Lincoln wrote the following note to his advisors. Use it to answer the question below. This morning, as for some days past, it seems exceedingly probable that this Administration will not be re-elected. Then it will be my duty to so cooperate with the President elect, as to save the Union between the election and the inauguration; as he will have secured his election on such ground that he can not possibly save it afterwards. inauguration:when a president takes office What did Abraham Lincoln think would happen if he lost the election? Choices: [(a) The election would bring the country back together., (b) The new president would not be able to save the Union., (c) The new president would ask Lincoln to be vice president.] Answer index: 1
@@
Subject: Social Studies Skill: mount-rushmore Description: When did workers start carving Mount Rushmore? Choices: [(a) in 1863, (b) in 1888, (c) in 1901, (d) in 1927] Answer index: 3
@@
Subject: Social Studies Skill: understand-overall-supply-and-demand Description: Eli Whitney invented the cotton gin in 1793. The cotton gin is a machine that makes it easier to harvest cotton quickly. What happened to the overall supply of cotton after 1793? Choices: [(a) The supply went up., (b) The supply went down.] Answer index: 0
Subject: Social Studies Skill: costs-and-benefits Description: Samuel is deciding whether to bake banana muffins or lemon muffins. He wants the muffins to be tasty. But he also wants to make a lot of muffins. Suppose Samuel decides to bake lemon muffins. Which result would be a cost? Choices: [(a) Samuel will get to eat more muffins. He can make more lemon muffins than banana muffins., (b) Samuel will give up the chance to eat banana muffins. He thinks banana muffins are tastier than lemon muffins.] Answer index: 1
@@
Subject: Social Studies Skill: antebellum-period-economies-of-the-north-and-south Description: In the United States, most enslaved people worked on large farms that grew crops to sell. What were these farms called? Choices: [(a) subsistence farms, (b) workshops, (c) factories, (d) plantations] Answer index: 3
@@
Subject: Social Studies Skill: jamestown-growth-of-a-colony Description: In 1613, a Jamestown settler kidnapped Chief Powhatan’s favorite daughter. What was she called? Choices: [(a) Pocahontas, (b) Massasoit, (c) Squanto, (d) Sacagawea] Answer index: 0
Subject: Social Studies Skill: world-war-ii-lead-up-to-war-in-europe-asia Description: After the Treaty of Versailles was signed, a German newspaper stated the following. Read the passage. Then answer the question. Today German honor is being carried to its grave. Do not forget it! The German people will, with unceasing labor, press forward to reconquer the place among the nations to which it is entitled. Then will come vengeance for the shame of 1919. vengeance:payback According to the passage, what does the author want Germany to do in the future? Choices: [(a) The author wants Germany to join the Allied powers., (b) The author wants Germany to become a powerful nation again., (c) The author wants Germany to forget about World War I., (d) The author wants Germany to join the League of Nations.] Answer index: 1
@@
Subject: Social Studies Skill: southern-colonies-founding-and-government Description: Many Virginians did not like the colonial government’s friendly policies toward Native Americans. In 1676, these Virginians violently rebelled against the government in Virginia. This event became known as Bacon’s Rebellion. Which of the following is an example of a violent rebellion? Choices: [(a) writing letters to the colonial governor of Virginia and the king, (b) voting in the next election for the Virginian colonial representatives, (c) marching on the colonial capital of Jamestown and burning it down] Answer index: 2
@@
Subject: Social Studies Skill: middle-colonies-economy-and-society Description: Merchants played a key role in trade between colonies and between the colonies and Europe. Which of these things would a merchant be likely to do? Choices: [(a) use lumber from the Middle Colonies to build houses, (b) grow cash crops such as corn and wheat, (c) sell corn from the Middle Colonies to people in England, (d) make iron products such as plows and horseshoes] Answer index: 2
@@
Subject: Social Studies Skill: world-war-ii-global-events Description: Kathryn Mary Doody was a nurse at an army hospital in Hawaii when Pearl Harbor was bombed. Read Kathryn Mary Doody’s account of the events of December 7, 1941. Then answer the question. Pretty soon the night nurse came off duty, and she said to us, "Girls, you know what’s happening?" And we said, "No, what’s happening?" She said, "The island of Oahu has been attacked by the enemy, Japan." And I said, "Shut up." Because she was a person that always had the greatest stories to relate. And she said, "Don’t pay any attention to me. Go turn your radios on." So I went in the room and turned the radio on and that’s exactly what they were repeating, over and over again.. . .So I thought, "That means war. What’s war going to be like?" According to the passage, what was Kathryn Mary Doody’s reaction when she first heard about the attack on Pearl Harbor? Choices: [(a) She was scared and went to church to pray for peace., (b) She didn’t believe that it was true., (c) She told all of her friends that they should help the military., (d) She was angry at the United States government.] Answer index: 1
Subject: Social Studies Skill: early-mesopotamia Description: Complete the text. Mesopotamia is an area between the ___ Rivers. Because of the rich soil created by the two rivers, Mesopotamia is often called the Fertile Crescent. Choices: [(a) Indus and Ganges, (b) Mississippi and Missouri, (c) Tigris and Euphrates, (d) Nile and Jordan] Answer index: 2
@@
Subject: Social Studies Skill: ancient-china Description: For people in ancient Chinese communities, flooding around the Yellow and Yangtze Rivers was a common problem. But floods had both positive and negative effects for people in ancient China. Which of the following was a positive effect of flooding? Choices: [(a) Floods killed many animals who lived near the river., (b) After a river flooded, fertile soil was left on the land., (c) Many crops were destroyed by the floods., (d) Floods destroyed homes located near the river.] Answer index: 1
Subject: Social Studies Skill: origins-of-islam Description: Muslims believe there is only one god. They call this god Allah, which means "God" in Arabic. The following passage describes a Muslim belief about how Allah revealed, or communicated, messages to his followers. Read the passage. Then answer the question below. We believe in Allah and [the message that] was revealed to us and what was revealed to Abraham . . . and the teachings which Allah gave to Moses and Jesus and to other prophets. Al Imran 3:84 in Zafar Ishaq Ansari, translator, Towards Understanding the Qur’an, Volume 1. Copyright 1988 by The Islamic Foundation. Based on this passage, which statement do most Muslims believe is true? Choices: [(a) Allah is more powerful than the god Jews and Christians worship., (b) Allah is the same god that Jews and Christians worship., (c) Allah sent messages to Abraham, but not to other prophets.] Answer index: 1
@@
Subject: Social Studies Skill: the-teachings-of-confucius Description: Confucius’s teachings have been passed down in a text known as The Analects (AN-el-ekts). Read the passage about The Analects. Then answer the question below. The Analects is made up of discussions between Confucius and his students. When Confucius was alive, his students all remembered their own conversations with Confucius. When Confucius died, his students gathered together to write down their memories of Confucius’s teachings. The text they wrote down was called The Analects, which means "the selected speeches." Complete the text. collected and wrote down The Analects. The Analects is made up of ___. Choices: [(a) Confucius . . . books Confucius wrote, (b) Confucius . . . conversations Confucius had with his students, (c) Confucius’s students . . . books Confucius wrote, (d) Confucius’s students . . . conversations Confucius had with his students] Answer index: 3
@@
Subject: Social Studies Skill: ancient-egypt-the-old-kingdom Description: Complete the text. Before Upper and Lower Egypt became one kingdom, they were divided into nomes, or ___. Each nome was controlled by a nomarch, or local king. Choices: [(a) territories, (b) farms, (c) churches, (d) teams] Answer index: 0
@@
Subject: Social Studies Skill: the-mauryan-empire Description: One of Chandragupta’s most important military and political advisors was a man named Kautilya. Read the following story about Kautilya and Chandragupta. Then answer the question below. Kautilya and Chandragupta gathered an army to overthrow the Nanda Empire. At first, their soldiers attacked the center of the Nanda Empire, but they were defeated by the powerful Nanda army. According to tradition, Kautilya and Chandragupta took that mother’s advice about eating hot bread and used it to defeat the Nanda Empire. How did Kautilya and Chandragupta use the mother’s advice to defeat the Nanda Empire? Choices: [(a) They convinced members of the Nanda army to change sides and help Chandragupta overthrow the empire., (b) They attacked the Nanda Empire from the edges, where it was weakest, rather than from the center where it was strong., (c) They attacked the center of the Nanda Empire until the army was too tired and weak to keep fighting.] Answer index: 1
Subject: Social Studies Skill: origins-of-christianity Description: Complete the text. Jesus of Nazareth is the most important figure in Christianity. Most Christians believe he ___. Choices: [(a) is God, (b) built a temple in Jerusalem, (c) was a famous military leader, (d) was a Roman emperor] Answer index: 0
@@
Subject: Social Studies Skill: ancient-egyptian-religion Description: In ancient Egypt, powerful gods were an important part of religious beliefs. People who believe in many gods are called polytheists (pol-ee-THEE-ists). Which statement would support the idea that ancient Egyptians were polytheists? Choices: [(a) Many ancient Egyptians worshipped the gods Osiris, Isis, and Horus., (b) Ancient Egyptian religious traditions lasted over 3,000 years., (c) Many different creation stories were popular in ancient Egypt., (d) Ancient Egypt was located in northeastern Africa, along the Nile River.] Answer index: 0
@@
Subject: Social Studies Skill: the-gupta-empire Description: The Gupta Empire was founded by a man who became known as Chandra Gupta (CHUN-druh GOOP-tuh). Read about Chandra Gupta. Then complete the sentence below. Chandra Gupta was born to a ruling family, but they only controlled a small area and were not very powerful. That began to change around 320 CE when Chandra Gupta married Princess Kumaradevi. Princess Kumaradevi came from a powerful family that controlled an area along the Ganges River. By marrying Kumaradevi, Chandra Gupta expanded the area he controlled, and he began forming an empire. Complete the following sentence. Chandra Gupta began the Gupta Empire after he ___. Choices: [(a) threatened to stop trading with regions that refused to let him rule, (b) inherited the large region his family had controlled for generations, (c) married a princess and took control of the area her family ruled, (d) won battles against many kingdoms along the Ganges River] Answer index: 2